Light bulbs are packaged in cartons of three. 10% of the bulbs are found to be faulty. Calculate the probability of finding two faulty bulbs in a single carton.
Answer by stanbon(75887)

Light bulbs are packaged in cartons of three. 10% of the bulbs are found to be faulty. Calculate the probability of finding two faulty bulbs in a single carton.
--------------
Binomial Problem with n = 3 and p = 0.10
---
P(x = 2) = 3C2*0.1^2*0.9 = 3*0.01*0.9 = 0.027
===============
Cheers,
Stan H.
=========